OpenStudy (anonymous):

The width of a rectangle is 17 inches. Its perimeter is 102 inches. Find the length of the rectangle

OpenStudy (nathan917):

Easy Question.

OpenStudy (nathan917):

Whats the Width.?

OpenStudy (anonymous):

17

OpenStudy (nathan917):

Ok, To find the length we must add the width on both sides.

OpenStudy (nathan917):

Both sides are the same.

OpenStudy (anonymous):

OMG...I JUST REALIZED IT WAS A RECTANGLE....NOT A TRIANGLE. SORRY I WAS THINKING OF THE FORMULA FOR THE TRIANGLE

OpenStudy (nathan917):

Lol.

OpenStudy (nathan917):

It's ok. :p

OpenStudy (anonymous):

17+17=34 102-34=68 68/2 = 34 Length= 34